OpenStudy (anonymous):

Eileen saves dimes and quarters. She has 40 coins which totals $6.55, how many of each coins does she has? (has to be a linear equation)

OpenStudy (anonymous):

(0.10)*(#dimes)+(0.25)*(#quarters)=6.55............................(eqn 1) (#dimes)+(#quarters)=40.....................................................(eqn 2)